OpenStudy (anonymous):

how to solve the exact 2nd order differential equation, (x^2)y''+xy'-y=0?

OpenStudy (amistre64):

there is a method of solving the characteristic quadratic ....

OpenStudy (amistre64):

or, power series are fun

OpenStudy (amistre64):

\[Let\\~\\y=\sum_0a_nx^n\]\[y'=\sum_0a_n~nx^{n-1}\]\[y''=\sum_0a_n~n(n-1)x^{n-2}\]

OpenStudy (amistre64):

forgot to up my indexes ... y' from 1, and y'' from 2 :)

OpenStudy (amistre64):

\[(x^2)y''+xy'-y=0\] \[(x^2)\sum_2a_n~n(n-1)x^{n-2}+x\sum_1a_n~nx^{n-1}-\sum_0a_nx^{n}=0\] \[\sum_2a_n~n(n-1)x^{n}+\sum_1a_n~nx^{n}+\sum_0-a_nx^{n}=0\] \[\sum_2a_n~n(n-1)x^{n}+\sum_1(a_n~n-a_n)x^{n}-a_1x=0\] \[\sum_2(a_n~n(n-1)+a_n~n-a_n)x^{n}+(a_1-a_1)x-a_1x=0\] \[\sum_2~a_n~(n(n-1)+n-1)x^{n}-a_1x=0\] this is identically zero when a1 = 0, and the recurrsion of the coeffs of the summation are zero